my grandmother age has a units digit of three, and the aquare of the tens digits is her age reversed. ho old is my grandmother?
Answer by stanbon(48558)

my grandmother age has a units digit of three, and the aquare of the tens digits is her age reversed. how old is my grandmother?
--------------
Assuming grandmother's age is a two digit number, it can be written
as 10t+3 where t is the ten's digit. The reverse is 3*10 + t
-----------
EQUATION:
t^2 = 3*10+t
t^2-t-30=0
(t-6)(t+5)=0
t=6
------------
Grandma is 63
==================
Cheers,
Stan H.